OpenStudy (anonymous):

What is the solution to the equation 5x = 40? 7 8 45 200

OpenStudy (anonymous):

i had gotten 8

OpenStudy (cgbork):

8

OpenStudy (lilg132):

8

OpenStudy (cgbork):

Divide each side by 5 to get x

pooja195 (pooja195):

right when you isolate x we divide 40/5=8 therefore x=8 :) good job

OpenStudy (lilg132):

5x = 40 5x / 5 = 40 / 5 x = 8

OpenStudy (anonymous):

thanks again lol :D